### v1

# Causal Effect Estimation under Networked Interference without Networked Unconfoundedness Assumption

Abstract
Estimating causal effects under networked interference is a crucial yet challenging problem. Existing methods based on observational data mainly rely on the networked unconfoundedness assumption, which guarantees the identification of networked effects. However, the networked unconfoundedness assumption is usually violated due to the latent confounders in observational data, hindering the identification of networked effects. Interestingly, in such networked settings, interactions between units provide valuable information for recovering latent confounders. In this paper, we identify three types of latent confounders in networked inference that hinder identification: those affecting only the individual, those affecting only neighbors, and those influencing both. Specifically, we devise a networked effect estimator based on identifiable representation learning techniques. Theoretically, we establish the identifiability of all latent confounders, and leveraging the identified latent confounders, we provide the networked effect identification result. Extensive experiments validate our theoretical results and demonstrate the effectiveness of the proposed method.
Machine Learning, ICML
1 Introduction
Estimating causal effects under network interference is a crucial yet challenging problem across various domains, including human ecology (Ferraro et al., 2019), advertising (Parshakov et al., 2020), and epidemiology (Barkley et al., 2020). The key challenge is that networked interference introduces interactions between units, violating the Stable Unit Treatment Value Assumption (SUTVA). For example, when evaluating the effect of a flu vaccine on individual infection rates, a standard causal inference approach assumes that an individual’s infection risk depends only on their own vaccination status. However, in reality, vaccination generates herd immunity effects—vaccinating one person may reduce disease transmission within the population, indirectly lowering the infection risk of others. This violation of SUTVA introduces bias into traditional causal inference methods, rendering standard estimands inapplicable (Forastiere et al., 2021). To model the interference between units, the existing methods focus on estimating three kinds of networked effects: main effects (effects of units’ own treatments), spillover effects (effects of units’ treatments on other units), and total effects (combined main and spillover effects).
Figure 1: A toy example showing networked interference between units. The networked interference introduces the interaction between units, i.e., the solid red arrows. Such arrows violate the traditional SUTVA assumption, leading to the non-identifiable problem.
To estimate causal effects from observational networked data, a series of works have been proposed under the networked unconfoundedness assumption. This assumption posits that no unobserved confounders exist beyond the observed covariates and the covariates of neighboring units. Under this assumption, Forastiere et al. (2021) establish the identification of network effects and propose the joint generalized propensity score for effect estimation. Building on this, Chin (2019); Ma & Tresp (2021); Cai et al. (2023) introduce balanced representation techniques to construct conditional outcome estimators for effect estimation. Additionally, Liu et al. (2019); Chen et al. (2024) develop doubly robust estimators to improve the robustness of network effect estimation under network interference.
Figure 2: Assumed causal graph in this paper. $x$ denotes observed proxies, $u$ denotes latent confounders, $t$ denotes the treatment, and $y$ denote the outcome of interest. We assume latent confounders $u$ contains three types of variables, i.e., $u^{i}$ affecting unit itself, $u^{n}$ affecting unit’s neighborhoods, and $u^{c}$ affecting both.
However, the networked unconfoundedness assumption is often violated in real-world scenarios, significantly limiting the effectiveness of existing methods. For example, in the case of flu vaccination, whether a person chooses to get vaccinated may depend on their income level or their family’s financial situation. However, such socioeconomic factors are often difficult to measure directly due to privacy concerns or data collection limitations. In such cases, latent confounders exist, violating the networked unconfoundedness assumption and introducing bias into existing methods.
To tackle the above challenge, we aim to develop a method that does not rely on the networked unconfoundedness assumption. Specifically, we begin by exploring three types of latent confounders, shown in Figure 2, which hinder the effect identification: $u^{i}$ affecting only the individual, $u^{n}$ affecting only neighbors, and $u^{c}$ influencing both. Rather than assuming networked unconfoundedness, we investigate the identifiability of latent confounders in the presence of networked interference. We found that networked interference provides additional auxiliary information that facilitates the identification of latent confounders. Built on the identified latent confounders, we theoretically establish the networked effect identification result and further devise an effect estimator under networked interference. Overall, our contribution can be summarized as follows:
  • We address the problem of networked effect identification and estimation in the presence of latent confounders. We categorize three types of latent confounders that hinder identification.
  • We explore the identifiability of latent confounders under networked interference, leveraging which, we achieve the networked effect identification.
  • We devise an estimator built on the theoretical findings. Extensive experiments validate our theoretical results and demonstrate the effectiveness of the proposed method.
2 Related Works
Classic Causal Inference has been studied in two languages: the graphical models (Pearl, 2009) and the potential outcome framework (Rubin, 1974). The most related method is the propensity score method in the potential outcome framework, e.g., IPW method (Rosenbaum & Rubin, 1983; Rosenbaum, 1987), which is widely applied to many scenarios (Rosenbaum & Rubin, 1985; Li et al., 2018; Cai et al., 2024). There are also many outcome regression models, including meta-learners (Künzel et al., 2019), neural networks-based works (Johansson et al., 2016; Assaad et al., 2021). By incorporating them, one can construct a doubly robust estimator (Robins et al., 1994), i.e., the effect estimator is consistent as either the propensity model or the outcome repression model is consistent.
Causal Inference without SUTVA has drawn increasing attention recently. Liu et al. (2016) extend the traditional propensity score to account for neighbors’ treatments and features and propose a generalized Inverse Probability Weighting (IPW) estimator. Forastiere et al. (2021) define the joint propensity score and then propose a subclassification-based method. Drawing upon previous works, Lee et al. (2021) consider two IPW estimators and derive a closed-form estimator for the asymptotic variance. Based on the representation learning, Ma & Tresp (2021) add neighborhood exposure and neighbors’ features as additional input variables and applies HSIC to learn balanced representations. Jiang & Sun (2022) use adversarial learning to learn balanced representations for better effect estimation. Ma et al. (2022) propose a framework to learn causal effects on a hypergraph. (Cai et al., 2023) propose a reweighted representation learning method to learn balanced representations. Under networked interference, McNealis et al. (2023); Liu et al. (2023); Chen et al. (2024) propose an estimator to achieve DR property. However, these works do assume the unconfoundedness assumption, which might not hold in real-world scenarios. Different from them, we explore the problem of networked effect estimation without the unconfoundedness assumption.
Causal Inference without Uncounfoundedness Assumption is an important problem since the unconfoundedness assumption is usually violated in observational studies. Classic methods to solve this problem usually assume there exist additional variables, e.g., instrumental variable (Pearl et al., 2000; Stock & Trebbi, 2003; Wu et al., 2022), proximal variable (Miao et al., 2018; Tchetgen Tchetgen et al., 2024). Another effective way to address this problem is to recover the latent confounder using representation learning methods. CEVAE (Louizos et al., 2017) assumes that latent confounders can be recovered by their proxies and applies VAE to learn confounders. As a follow-up work, TEDVAE (Zhang et al., 2021) decouples the learned latent confounders into several factors to achieve a more accurate estimation of treatment effects. In the mediation analysis, DMAVAE (Xu et al., 2023) proposes to recover latent confounders using the VAE similar to CEVAE. Our work is closely related to these works. Different from them, we focus on the causal effect without the unconfoundedness assumption in the presence of networked interference. We also provide theoretical guarantees for the latent confounder identifiability, which ensures the effectiveness of our estimator.
3 Notations, Assumptions, Esitimands
In this section, we start with the notations used in this work. Let $U\in\mathcal{U}$ be the latent confounders and also let $X\in\mathcal{X}$ be the proxies of latent confounders. Let $T\in\{0,1\}$ denote a binary treatment, where $T=1$ indicates a unit receives the treatment (treated) and $T=0$ indicates a unit receives no treatment (control). Let $Y\in\mathcal{Y}$ be the outcome. We assume that $U$ can be decomposed to $U^{i}$ affecting the unit itself, $U^{c}$ affecting the neighborhood, and $U^{n}$ affecting both. Let lowercase letters (e.g., $x,y,t$) denote the value of random variables. Let lowercase letters with subscript $i$ denote the value of the specified $i$-th unit. Thus, a network dataset is denoted as $D=(\{x_{i},t_{i},y_{i}\}_{i=1}^{n},E)$, where $E$ denotes the adjacency matrix of network and $n$ is the total number of units. We denote the set of first-order neighbors of $i$ as $\mathcal{N}_{i}$ and denote the treatment and feature vectors received by unit $i$’s neighbors as $t_{\mathcal{N}_{i}}$ and $x_{\mathcal{N}_{i}}$. Due to the presence of networked interference, a unit’s potential outcome is influenced not only by its treatment but also by its neighbors’ treatments, and thus the potential outcome is denoted by $y_{i}(t_{i},t_{\mathcal{N}_{i}})$. The observed outcome $y_{i}$ is known as the factual outcome, and the remaining potential outcomes are known as counterfactual outcomes.
Further, following Forastiere et al. (2021); Chen et al. (2024), we assume that the dependence between the potential outcome and the neighbors’ treatments is through a specified summary function $g$: $\{0,1\}^{|\mathcal{N}_{i}|}\rightarrow[0,1]$, and let $z_{i}$ be the neighborhood exposure given by the summary function, i.e., $z_{i}=g(t_{\mathcal{N}_{i}})$. We aggregate the information of the neighbors’ treatments to obtain the neighborhood exposure by $z_{i}=\frac{\sum_{j\in\mathcal{N}_{i}}t_{j}}{|\mathcal{N}_{i}|}$. Therefore, the potential outcome $y_{i}(t_{i},t_{\mathcal{N}_{i}})$ can be denoted as $y_{i}(t_{i},z_{i})$, which means that under networked interference, each unit is affected by two kinds of treatments: the binary individual treatment $t_{i}$ and the continuous neighborhood exposure $z_{i}$.
In this paper, our goal is to estimate the average dose-response function, as well as the conditional average dose-response function:
$\psi(t,z):=\mathbb{E}[Y(t,z)],$ (1)
$\mu(t,z,x,x_{\mathcal{N}}):=\mathbb{E}[Y(t,z)|X=x,X_{\mathcal{N}}=x_{\mathcal{N}}],$
Based on the average dose-response function, existing works mostly focus on the following causal effects:
Definition 3.1 (Average Main Effects (AME) ).
AME measures the difference in mean outcomes between units assigned to $T=t,Z=0$ and assigned $T=t^{\prime},Z=0$: $\tau^{(t,0),(t^{\prime},0)}=\psi(t,0)-\psi(t^{\prime},0)$.
Definition 3.2 (Average Spillover Effects (ASE) ).
ASE measures the difference in mean outcomes between units assigned to $T=0,Z=z$ and assigned $T=0,Z=z^{\prime}$: $\tau^{(0,z),(0,z^{\prime})}=\psi(0,z)-\psi(0,z^{\prime})$.
Definition 3.3 (Average Total Effects (ATE) ).
ATE measures the difference in mean outcomes between units assigned to $T=t,Z=z$ and assigned $T=t^{\prime},Z=z^{\prime}$: $\tau^{(t,z),(t^{\prime},z^{\prime})}=\psi(t,z)-\psi(t^{\prime},z^{\prime})$.
Definition 3.4 (Individual Main Effects (IME) ).
IME measures the difference in mean outcomes of a particular unit $x_{i}$ assigned to $T=t,Z=0$ and assigned $T=t^{\prime},Z=0$: $\tau_{i}(x_{i},x_{\mathcal{N}_{i}})^{(t,0),(t^{\prime},0)}=\mu(x_{i},x_{\mathcal{N}_{i}},t,0)-\mu(x_{i},x_{\mathcal{N}_{i}},t^{\prime},0)$.
Definition 3.5 (Individual Spillover Effects (ISE) ).
ISE measures the difference in mean outcomes of a particular unit $x_{i}$ assigned to $T=0,Z=z$ and assigned $T=0,Z=z^{\prime}$: $\tau_{i}(x_{i},x_{\mathcal{N}_{i}})^{(0,z),(0,z^{\prime})}=\mu(x_{i},x_{\mathcal{N}_{i}},0,z)-\mu(x_{i},x_{\mathcal{N}_{i}},0,z^{\prime})$.
Definition 3.6 (Individual Total Effects (ITE) ).
ITE measures the difference in mean outcomes of a particular unit $x_{i}$ assigned to $T=t,Z=z$ and assigned $T=t^{\prime},Z=z^{\prime}$: $\tau_{i}(x_{i},x_{\mathcal{N}_{i}})^{(t,z),(t^{\prime},z^{\prime})}=\mu(x_{i},x_{\mathcal{N}_{i}},t,z)-\mu(x_{i},x_{\mathcal{N}_{i}},t^{\prime},z^{\prime})$.
The main effects reflect the effects of changing neighborhood exposure $t$ to $t^{\prime}$. The spillover effects reflect the effects of changing neighborhood exposure $z$ to $z^{\prime}$. And the total effects represent the combined effect of both main effects and spillover effects.
Throughout this paper, we also assume the following assumptions hold:
Assumption 3.7 (Network Consistency).
The potential outcome is the same as the observed outcome under the same individual treatment and neighborhood exposure, i.e., $y_{i}=y_{i}(t_{i},z_{i})$ if unit $i$ actually receives $t_{i}$ and $z_{i}$.
Assumption 3.8 (Network Overlap).
Given any individual and neighbors’ features, any treatment pair $(t,z)$ has a non-zero probability of being observed in the data, i.e., $\forall x_{i},x_{\mathcal{N}_{i}},t_{i},z_{i},\quad 0<p(t_{i},z_{i}|x_{i},x_{\mathcal{N}_{i}})<1$.
Assumption 3.9 (Neighborhood Interference).
The potential outcome of a unit is only affected by their own and the first-order neighbors’ treatments, and the effect of the neighbors’ treatments is through a summary function: $g$, i.e., $\forall t_{\mathcal{N}_{i}}$,$t^{\prime}_{\mathcal{N}_{i}}$ which satisfy $g(t_{\mathcal{N}_{i}})=g(t^{\prime}_{\mathcal{N}_{i}})$, the following equation holds: $y_{i}(t_{i},t_{\mathcal{N}_{i}})=y_{i}(t_{i},t^{\prime}_{\mathcal{N}_{i}})$.
These assumptions are commonly assumed in existing causal inference methods such as Forastiere et al. (2021); Cai et al. (2023); Ma et al. (2022). Specifically, Assumption 3.7 states that there can not be multiple versions of a treatment. Assumption 3.8 requires that the treatment assignment is nondeterministic. Assumption 3.9 rules out the dependence of the outcome of unit $i$, $y_{i}$, from the treatment received by units outside its neighborhood, i.e., $t_{j},j\notin\mathcal{N}_{i}$, but allows $y_{i}$ to depend on the treatment received by his neighbors, i.e., $t_{k},k\in\mathcal{N}_{i}$. Also, Assumption 3.9 states the interaction dependence is assumed to be through a summary function $g$. Note that Assumption 3.9 is reasonable in reality for some reason. First, in many applications units are affected by their first-order neighbors, and the affection of higher-order neighbors is also transported through the first-order neighbors. Second, it is also reasonable that a unit is affected by a specific function of other units’ treatment, e.g., how much job-seeking pressure a unit has will depend on how many of its friends receive job training.
Existing methods additionally assume the following assumption:
Assumption 3.10 (Networked Unconfoundedness).
The individual treatment and neighborhood exposure are independent of the potential outcome given the individual and neighbors’ features, i.e., $\forall t,z,\quad y_{i}(t,z){\perp\!\!\!\perp}t_{i},z_{i}|x_{i},x_{\mathcal{N}_{i}}$.
Assumption 3.10 is an extension of the traditional unconfoundedness assumption and indicates that there is no unmeasured confounder which is the common cause of $y_{i}$ and $t_{i},z_{i}$.
Under the assumptions above, the networked effects can be identified (Forastiere et al., 2021; Cai et al., 2023). However, Assumption 3.10 might be too strong to hold, since we can not promise that all confounders are observed in real-world scenarios. Instead, we assume a much weaker assumption by incorporating the latent variables as follows:
Assumption 3.11 (Latent Networked Unconfoundedness).
The individual treatment and neighborhood exposure are independent of the potential outcome given the latent individual and neighbors’ confounders, i.e., $\forall t,z,\quad y_{i}(t,z){\perp\!\!\!\perp}t_{i},z_{i}|u^{i}_{i},u^{c}_{i},u^{c}_{\mathcal{N}_{i}},u^{n}_{\mathcal{N}_{i}}$.
This assumption allows for the latent confounders $u^{i},u^{c},u^{n}$. Here, we recognize three types of latent confounders. What serves as an adjustment set under the networked setting is the units’ $u^{i},u^{c}$ and the neighbors’ $u^{c}_{\mathcal{N}_{i}},u^{n}_{\mathcal{N}_{i}}$. This also motivates us to identify each latent confounder for a better estimation. In the next section, we will introduce the identifiability of each latent confounder, and further achieve networked effect identification.
4 Networked Causal Effect Identification via Representation Learning
To begin with, following existing identifiable representation learning methods (Khemakhem et al., 2020; Lu et al., 2022), we first introduce the generative model as follows:
$p_{\bm{\theta}}(X,U|X_{\mathcal{N}})$ $=p_{\bm{f}}(X|U)p_{\bm{T},\bm{\lambda}}(U|X_{\mathcal{N}})$ (2)
$p_{\bm{f}}(X|U)$ $=p_{\bm{\epsilon}}(X-\bm{f}(U))$
Further, we assume $p_{\bm{T},\bm{\lambda}}(U|X_{\mathcal{N}})$ follows the exponential family distribution.
Assumption 4.1.
The correlation between $U$ and $X_{\mathcal{N}}$ is characterized by:
$p_{\bm{T},\bm{\lambda}}(U|X_{\mathcal{N}})$ $=\frac{\mathcal{Q}(U)}{\mathcal{C}(X_{\mathcal{N}})}\exp\left[\bm{T}(U)^{T}\bm{\lambda}(X_{\mathcal{N}})\right]$ (3)
where $\mathcal{Q}$ is the base measure, $\mathcal{C}$ is the normalizing constant. The $\bm{\lambda}(X_{\mathcal{N}})$ is an arbitrary function, and the sufficient statistics $\bm{T}(U)=[\bm{T}_{f}(U)^{T},\bm{T}_{MLP}(U)^{T}]^{T}$ contains a) the sufficient statistics $\bm{T}_{f}(U)^{T}=[\bm{T}_{1}(U_{(1)})^{T},\dots,\bm{T}_{1}(U_{(d_{U})})^{T}]$ of a factorized exponential family, where all the $\bm{T}_{i}(U_{(i)})$ have dimension larger or equal to $2$ and $d_{U}$ is the dimension of $U$, and b) the output $\bm{T}_{MLP}(U)$ of a neural network with ReLU activations.
This assumption is introduced by Lu et al. (2022). The distribution in Assumption 4.1 is more flexible than the standard assumed distribution condition in identifiable representation learning (Eq. (7) in Khemakhem et al. (2020)). This assumption allows for the case that the different elements of latent confounders are not independent given the conditional set. The term $\bm{T}_{MLP}(U)$ does capture arbitrary dependencies between latent variables since the neural network with ReLU activation has universal approximation power.
Now, we formally state the theoretical result of the identifiability of latent confounders.
Theorem 4.2.
Suppose Assumption 4.1 holds, and suppose the following conditions hold: (1) The set $\{X\in\mathcal{O}|\varphi_{\bm{\epsilon}}(X)=0\}$ has measure zero where $\varphi_{\bm{\epsilon}}$ is the characteristic function of density $p_{\bm{\epsilon}}$. (2) $\bm{f}$ is injective and has all second-order cross derivatives. (3) The sufficient statistics in $\bm{T}_{\bm{f}}$ are all twice differentiable. (4) There exist $k+1$ distinct values $x_{\mathcal{N}_{0}},...,x_{\mathcal{N}_{k+1}}$ such that the matrix
$L=($ $\bm{\lambda}(x_{\mathcal{N}_{1}})-\bm{\lambda}(x_{\mathcal{N}_{0}}),$
$\bm{\lambda}(x_{\mathcal{N}_{k+1}})-\bm{\lambda}(x_{\mathcal{N}_{0}}))$
of size $k\times k$ is invertible where $k=|U^{i}|+|U^{c}|+|U^{n}|$ is the dimension of latent variables. Then we learn the true latent variable $U^{i},U^{c},U^{n}$ up to a permutation and simple transformations.
Discussion on assumptions and conditions. Assumption 4.1 indicates our theory holds for a rich family of conditional densities (Wainwright & Jordan, 2008). The assumption on the exponential family distribution is not strong, since many well-known distributions belong to this family, including Gaussian, Uniform, Poisson distributions, and so on. Condition (1)-(4) is a common assumption in representation learning in causal representation learning, e.g., (Khemakhem et al., 2020; Lu et al., 2022). Notably, the most important condition is the condition (4) which requires that the auxiliary information should be sufficient enough. Under our networked setting, it requires that there exist enough distinct values of neighbors’ covariates, which is easy to hold if we collect enough covariates, especially when some of the covariates are continuous.
Theorem 4.2 indicates that, under mild assumptions, the latent confounders can be recovered up to a simple function, i.e., the recovered $\hat{U}^{i},\hat{U}^{c},\hat{U}^{n}$ satisfying $\hat{U}^{i}=h_{i}(U^{i}),\hat{U}^{c}=h_{c}(U^{c}),\hat{U}^{n}=h_{n}(U^{n})$ for some simple functions $h_{i}.h_{c},h_{n}$. Based on Theorem 4.2 above, we can further identify networked effect as follows:
Theorem 4.3.
Suppose Assumption 3.7, 3.9, 3.8, 3.11, and Theorem 4.2 holds, the networked effects $\psi(t,z)$ and $\mu(t,z,x,x_{\mathcal{N}})$ are identifiable.
Theorem 4.3 indicates that if we can identify latent confounders, the networked effect is thereby identifiable. This necessitates the utilization of identifiable representation learning techniques for causal inference in the presence of networked interference and latent confounders.
Figure 3: Model architecture of our proposed method. The representation learning module aims to learn the unobserved confounders. The feature module aggregates the information of covariates of unit $i$ and its neighbor. The outcome estimator module aims to estimate potential outcomes of unit $i$.
5 Methodology
In this section, leveraging the theoretical findings, we devise our networked effect estimator in the presence of latent confounders. Specifically, our estimator contains three modules, including the representation learning module, the feature module, and the outcome estimator module. The representation learning module is built on Theorem 4.2, aiming to correctly recover three types of latent confounders $u^{i},u^{c},u^{n}$. The feature module is built on Theorem 4.3, aggregating the information from units’ and neighbors’ information. This aggregated information is then input into the outcome estimator module to predict the networked causal effects. Overall, our model architecture is shown in Figure 3.
5.1 Representation Learning Module
Following existing work (Guo et al., 2020; Ma & Tresp, 2021; Jiang & Sun, 2022; Chen et al., 2024), we use Graph Convolution Networks (GCN (Defferrard et al., 2016; Kipf & Welling, 2016)) to aggregate the information of covariates of unit $i$ and its neighbors, i.e., $x_{i},x_{\mathcal{N}_{i}}$:
$h^{neigh}_{i,1}=\sigma(\sum_{j\in\mathcal{N}_{i}}\frac{1}{\sqrt{d_{i}d_{j}}}W_{1}^{T}x_{j}),$
$h_{i,2}=MLP_{1}(h^{neigh}_{i,1},x_{i}),$
where $\sigma(\cdot)$ is a non-linear activation function, $d_{i}$ is the degrees of unit $i$, $W_{1}$ is the learning weight matrix of GCN, and $MLP_{1}$ is a multilayer perception (MLP).
Then, given $x_{i}$ and $x_{\mathcal{N}_{i}}$, we employ the identifiable representation learning technique (Lu et al., 2022) to recover latent confounders. Specifically, we parametrize the prior following Assumption 4.1:
$p(u_{i}|x_{\mathcal{N}_{i}})$ (4)
$=$ $\langle MLP_{2}(u_{i}),MLP_{3}(h_{i,2})\rangle+\langle[u_{i},u_{i}^{2}],MLP_{4}(h_{i,2})\rangle$
where $u_{i}=[u_{i}^{i},u_{i}^{c},u_{i}^{n}]$ and $MLP_{2}(u_{i})$ serves as $\bm{T}_{MLP}$, the concatenated $[u_{i},u_{i}^{2}]$ serves as $\bm{T}_{f}$, $MLP_{3}(h_{i,2})$ serves as $\bm{\lambda}_{MLP}$, and $MLP_{4}(h_{i,2})$ serves as $\bm{\lambda}_{f}$ in Assumption 4.1.
As for the encoder, the variational approximation of the posterior is defined as:
$q(u^{i}_{i},u^{c}_{i},u^{n}_{i}|x_{i},x_{\mathcal{N}_{i}})=\prod_{i=0}^{|U|}\mathcal{N}(\mu=\hat{\mu}_{u_{i}},\sigma^{2}=\hat{\sigma}_{u_{i}}^{2}),$ (5)
where $\hat{\mu}_{u_{i}}$ and $\hat{\sigma}_{u_{i}}^{2}$ are the mean and variance of the Gaussian distribution parametrized by MLPs using $h_{i,2}$ as input.
As for the decoder, for a continuous outcome, we parametrize the probability distribution as a Gaussian distribution with its mean given by an MLP and a fixed variance $v^{2}$. For a discrete outcome, we use a Bernoulli distribution parametrized by an MLP similarly:
$p(x_{i}|u^{i}_{i},u^{c}_{i},u^{n}_{i})=\prod_{i=0}^{|X|}\mathcal{N}(\mu=\hat{\mu}_{x},\sigma^{2}=v_{x}^{2})$ (6)
or $p(x_{i}|u^{i}_{i},u^{c}_{i},u^{n}_{i})=\prod_{i=0}^{|X|}\mathbf{Bern}(\pi=\hat{\pi}_{x}),$
where for the continuous case $\hat{\mu}_{x}$ is the mean of the Gaussian distribution parametrized by an MLP using the sampled $u^{i}_{i},u^{c}_{i},u^{n}_{i}$ from posterior as input, and $v_{x}^{2}$ is the fixed variance of Gaussian distribution, and for the discrete case $\hat{\pi}_{x}$ is the mean of Bernoulli distribution similarly parametrized by an MLP.
For this module, we use the negative variational Evidence Lower BOund (ELBO) as the loss function, defined as
$\textbf{ELBO}={\mathbb{E}}_{q(u^{i},u^{c},u^{n}|x.x_{\mathcal{N}})}[\log p(x|u^{i},u^{c},u^{n})$ (7)
$+\log p(u^{i},u^{c},u^{n}|x_{\mathcal{N}})-\log q(u^{i},u^{c},u^{n}|x.x_{\mathcal{N}})].$
Following Lu et al. (2022), we utilize the score matching technique (Vincent, 2011) for training unnormalized probabilistic models to learn the parameters of $\bm{T}$ and $\bm{\lambda}$ by minimizing
$\mathcal{L}_{sm}={\mathbb{E}}_{q(u^{i},u^{c},u^{n}|x.x_{\mathcal{N}})}[$ (8)
$\|\nabla_{u}\log q(u^{i},u^{c},u^{n}|x.x_{\mathcal{N}})-\nabla_{u}\log p(u^{i},u^{c},u^{n}|x_{\mathcal{N}})\|^{2}],$
where $\nabla$ is the gradient operator.
5.2 Feature Module and Outcome Estimator Module
After obtaining $u^{i},u^{c},u^{n}$, we can aggregate the necessary information for the effect estimation. Specifically, in the feature module, we first aggregate the neighbors’ $u^{c},u^{n}$ to obtain $u^{c}_{\mathcal{N}},u^{n}_{\mathcal{N}}$ via GCN:
$h_{i,3}^{neigh}=\sigma(\sum_{j\in\mathcal{N}_{i}}\frac{1}{\sqrt{d_{i}d_{j}}}W_{2}^{T}[u^{c}_{j},u^{n}_{j}]),$
$h_{i,4}=MLP_{1}(h^{neigh}_{i,3},u^{i}_{i},u^{c}_{i}),$
where $\sigma(\cdot)$ is a non-linear activation function, $d_{i}$ is the degrees of unit $i$, $W_{2}$ is the learning weight matrix of GCN.
Then, we use $h_{i,4}$, $u^{i}_{i},u^{c}_{i}$ and $t_{i},z_{i}$ together to estimate $y_{i}$ of treated and control groups respectively, i.e.,
$\small\mu^{NN}(t_{i},z_{i},u^{i}_{i},u^{c}_{i},u^{c}_{\mathcal{N}_{i}},u^{n}_{\mathcal{N}_{i}})=\left\{\begin{aligned} MLP_{3}(z_{i},h_{i,2})&\quad t_{i}=1\\ MLP_{4}(z_{i},h_{i,2})&\quad t_{i}=0\end{aligned}\right.$
and the loss function is
$\mathcal{L}_{y}=\Sigma_{i=1}^{n}(y_{i}-\mu^{NN}(t_{i},z_{i},x_{i},x_{\mathcal{N}_{i}}))^{2}.$ (9)
Moreover, inspired by Jiang & Sun (2022); Cai et al. (2023), we further consider a balancing regularization term as our loss:
$\mathcal{L}_{IPM}=\text{IPM}(p(h_{i,4},t_{i},z_{i}),p(h_{i,4})p(t_{i})p(z_{i}))),$ (10)
where $\text{IPM}(p.q)=\sup_{g\in\mathcal{G}}|\int_{\mathcal{X}}g(x)(p(x)-q(x))dx|$ is the integral probability metric, measuring the distance between two distribution $p,q$, which can be implemented by Wasserstein Distance. Here the samples from $p(h_{i,4})p(t_{i})p(z_{i}))$ is obtained by randomly permuting $t_{i}$ and $z_{i}$ separately.
Overall, our final loss function is
$\mathcal{L}_{all}=-\textbf{ELBO}+\mathcal{L}_{sm}+\mathcal{L}_{y}+\mathcal{L}_{IPM}.$ (11)
6 Experiments
In this section, we validate the proposed method on two commonly used semisynthetic datasets. In detail, we verify the effectiveness of our algorithm and further evaluate the correctness of the analysis with the help of semisynthetic datasets.
Table 1: Experimental results on Flickr(homo) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
$\varepsilon_{average}$ $\varepsilon_{individual}$
Within Sample Out-of Sample Within Sample Out-of Sample
Methods AME ASE ATE AME ASE ATE IME ISE ITE IME ISE ITE
TARNET+z $0.0783_{\pm 0.0418}$ $0.0874_{\pm 0.0213}$ $0.2025_{\pm 0.0396}$ $0.0976_{\pm 0.0506}$ $0.0724_{\pm 0.0184}$ $0.1356_{\pm 0.0587}$ $0.1362_{\pm 0.0254}$ $0.1103_{\pm 0.0194}$ $0.2358_{\pm 0.0383}$ $1.0869_{\pm 1.2258}$ $0.1011_{\pm 0.0185}$ $1.0889_{\pm 1.2270}$
CFR+z $0.0579_{\pm 0.0247}$ $0.0785_{\pm 0.0070}$ $0.1651_{\pm 0.0121}$ $0.0507_{\pm 0.0192}$ $0.0783_{\pm 0.0070}$ $0.1581_{\pm 0.0097}$ $0.0599_{\pm 0.0240}$ $0.0786_{\pm 0.0070}$ $0.1654_{\pm 0.0120}$ $0.3465_{\pm 0.4615}$ $0.0786_{\pm 0.0069}$ $0.4102_{\pm 0.4278}$
GEst $0.1551_{\pm 0.0130}$ $0.2475_{\pm 0.0476}$ $0.0805_{\pm 0.0325}$ $0.1511_{\pm 0.0137}$ $0.2494_{\pm 0.0470}$ $0.0805_{\pm 0.0278}$ $0.1779_{\pm 0.0122}$ $0.2656_{\pm 0.0378}$ $0.1268_{\pm 0.0160}$ $0.2867_{\pm 0.2172}$ $0.2677_{\pm 0.0372}$ $0.2471_{\pm 0.2352}$
ND+z $0.1416_{\pm 0.0240}$ $0.0204_{\pm 0.0093}$ $0.0478_{\pm 0.0216}$ $0.1435_{\pm 0.0364}$ $0.0226_{\pm 0.0101}$ $0.0485_{\pm 0.0236}$ $0.1427_{\pm 0.0246}$ $0.0221_{\pm 0.0090}$ $0.0501_{\pm 0.0178}$ $0.3849_{\pm 0.2395}$ $0.0348_{\pm 0.0078}$ $0.3453_{\pm 0.2772}$
NetEst $0.0515_{\pm 0.0538}$ $0.0355_{\pm 0.0317}$ $0.0715_{\pm 0.0381}$ $0.0470_{\pm 0.0500}$ $0.0338_{\pm 0.0330}$ $0.0529_{\pm 0.0395}$ $0.0844_{\pm 0.0406}$ $0.0566_{\pm 0.0253}$ $0.1043_{\pm 0.0312}$ $0.2934_{\pm 0.3001}$ $0.2809_{\pm 0.3387}$ $0.3068_{\pm 0.1860}$
TNet $0.0319_{\pm 0.0249}$ $0.0274_{\pm 0.0309}$ $0.0735_{\pm 0.0240}$ $0.0299_{\pm 0.0231}$ $0.0277_{\pm 0.0313}$ $0.0715_{\pm 0.0214}$ $0.0347_{\pm 0.0282}$ $0.0276_{\pm 0.0313}$ $0.0752_{\pm 0.0263}$ $0.0561_{\pm 0.0648}$ $0.0286_{\pm 0.0331}$ $0.0918_{\pm 0.0555}$
Ours_w/o_IPM $0.0359_{\pm 0.0262}$ $\textbf{0.0133}_{\pm 0.0050}$ $0.0598_{\pm 0.0366}$ $0.0394_{\pm 0.0262}$ $\textbf{0.0123}_{\pm 0.0060}$ $0.0602_{\pm 0.0355}$ $0.0410_{\pm 0.0228}$ $\textbf{0.0155}_{\pm 0.0033}$ $0.0643_{\pm 0.0337}$ $0.0424_{\pm 0.0242}$ $\textbf{0.0142}_{\pm 0.0044}$ $0.0632_{\pm 0.0334}$
Ours $\textbf{0.0296}_{\pm 0.0219}$ $0.0252_{\pm 0.0208}$ $\textbf{0.0266}_{\pm 0.0208}$ $\textbf{0.0289}_{\pm 0.0208}$ $0.0252_{\pm 0.0208}$ $\textbf{0.0260}_{\pm 0.0200}$ $\textbf{0.0297}_{\pm 0.0220}$ $0.0252_{\pm 0.0208}$ $\textbf{0.0267}_{\pm 0.0209}$ $\textbf{0.0289}_{\pm 0.0209}$ $0.0252_{\pm 0.0208}$ $\textbf{0.0261}_{\pm 0.0201}$
6.1 Experimental Setup
Table 2: Experimental results on Flickr(hete) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
$\varepsilon_{average}$ $\varepsilon_{individual}$
Within Sample Out-of Sample Within Sample Out-of Sample
Methods AME ASE ATE AME ASE ATE IME ISE ITE IME ISE ITE
TARNET+z $0.1315_{\pm 0.0740}$ $0.1673_{\pm 0.0423}$ $0.3590_{\pm 0.0785}$ $0.1554_{\pm 0.1110}$ $0.1319_{\pm 0.0307}$ $0.2728_{\pm 0.1321}$ $0.2320_{\pm 0.0432}$ $0.2042_{\pm 0.0479}$ $0.4254_{\pm 0.0802}$ $1.5274_{\pm 1.6256}$ $0.1760_{\pm 0.0351}$ $1.5957_{\pm 1.5779}$
CFR+z $0.1131_{\pm 0.0476}$ $0.1437_{\pm 0.0081}$ $0.2960_{\pm 0.0182}$ $0.0998_{\pm 0.0458}$ $0.1412_{\pm 0.0085}$ $0.2789_{\pm 0.0309}$ $0.1445_{\pm 0.0407}$ $0.1463_{\pm 0.0083}$ $0.3242_{\pm 0.0224}$ $0.5946_{\pm 0.7379}$ $0.1448_{\pm 0.0087}$ $0.7104_{\pm 0.6761}$
GEst $0.3283_{\pm 0.0426}$ $0.4717_{\pm 0.1336}$ $0.1074_{\pm 0.0255}$ $0.3356_{\pm 0.0270}$ $0.4723_{\pm 0.1312}$ $0.0969_{\pm 0.0099}$ $0.3697_{\pm 0.0386}$ $0.5123_{\pm 0.1231}$ $0.2178_{\pm 0.0214}$ $0.7124_{\pm 0.6463}$ $0.5144_{\pm 0.1202}$ $0.5914_{\pm 0.7073}$
ND+z $0.2420_{\pm 0.0330}$ $0.0293_{\pm 0.0113}$ $0.0852_{\pm 0.0365}$ $0.2433_{\pm 0.0539}$ $0.0318_{\pm 0.0134}$ $0.0785_{\pm 0.0422}$ $0.2571_{\pm 0.0348}$ $0.0430_{\pm 0.0040}$ $0.1607_{\pm 0.0188}$ $0.5156_{\pm 0.2180}$ $0.0669_{\pm 0.0059}$ $0.4720_{\pm 0.2580}$
NetEst $0.0530_{\pm 0.0423}$ $0.0452_{\pm 0.0351}$ $0.0723_{\pm 0.0319}$ $0.0466_{\pm 0.0322}$ $0.0565_{\pm 0.0454}$ $0.0818_{\pm 0.0379}$ $0.1145_{\pm 0.0278}$ $0.0667_{\pm 0.0267}$ $0.1660_{\pm 0.0163}$ $0.6855_{\pm 0.2607}$ $0.5353_{\pm 0.2507}$ $0.5625_{\pm 0.1367}$
TNet $0.0411_{\pm 0.0238}$ $0.0206_{\pm 0.0073}$ $0.0282_{\pm 0.0297}$ $0.0417_{\pm 0.0237}$ $0.0196_{\pm 0.0098}$ $0.0268_{\pm 0.0314}$ $0.0936_{\pm 0.0170}$ $0.0338_{\pm 0.0065}$ $0.1360_{\pm 0.0210}$ $0.0950_{\pm 0.0157}$ $0.0361_{\pm 0.0074}$ $0.1415_{\pm 0.0223}$
Ours_w/o_IPM $\textbf{0.0296}_{\pm 0.0248}$ $\textbf{0.0125}_{\pm 0.0131}$ $0.0581_{\pm 0.0321}$ $\textbf{0.0254}_{\pm 0.0212}$ $\textbf{0.0171}_{\pm 0.0195}$ $0.0437_{\pm 0.0241}$ $0.0929_{\pm 0.0181}$ $0.0361_{\pm 0.0114}$ $0.1503_{\pm 0.0221}$ $\textbf{0.0910}_{\pm 0.0153}$ $0.0393_{\pm 0.0149}$ $0.1480_{\pm 0.0172}$
Ours $0.0377_{\pm 0.0207}$ $0.0206_{\pm 0.0157}$ $\textbf{0.0228}_{\pm 0.0058}$ $0.0384_{\pm 0.0225}$ $0.0226_{\pm 0.0177}$ $\textbf{0.0197}_{\pm 0.0088}$ $\textbf{0.0922}_{\pm 0.0115}$ $\textbf{0.0353}_{\pm 0.0121}$ $\textbf{0.1326}_{\pm 0.0161}$ $0.0941_{\pm 0.0100}$ $\textbf{0.0390}_{\pm 0.0135}$ $\textbf{0.1379}_{\pm 0.0164}$
Datasets We consider two wildly used semisynthetic datasets BlogCatalog and Flickr to verify the effectiveness of our estimator. We further use a synthetic dataset to validate the correctness of our theories, i.e., whether our method can correctly recover latent confounders.
Following existing works (Jiang & Sun, 2022; Guo et al., 2020; Ma et al., 2021; Chen et al., 2024), we use two semisynthetic datasets to evaluate our proposed method:
  • BlogCatalog (BC) is an online community where users post blogs. In this dataset, each unit is a blogger and each edge is the social link between units. The features are bag-of-words representations of keywords in bloggers’ descriptions.
  • Flickr is an online social network where users can share images and videos. In this dataset, each unit is a user and each edge is the social relationship between units. The features are the list of tags of units’ interests.
We reuse the original covariates as the latent confounders and then divide them into $u^{i},u^{c},u^{n}$. We generate the proxies $x$ using $x_{i}=w_{1}u_{i}+e_{i}$ where $w_{1}$ are randomly sampled from Uniform distribution $\mathcal{U}(0.5,1)$ and $e_{x,i}$ is standard Gaussian noise. Then given the latent confounder $u^{i}_{i},u^{c}_{i},u^{n}_{i}$ of unit $i$, the treatments are simulated by
$t_{i}=\begin{cases}1&\text{if \quad$tpt_{i}>\overline{tpt}$},\\ 0&\text{else},\end{cases}$
where $\overline{tpt}$ is the average of all $tpt_{i}$, and $tpt_{i}=pt_{i}+pt_{\mathcal{N}_{i}}$, and $pt_{i}=Sigmoid(w_{2}\times[u^{i}_{i},u^{c}_{i}])$, and $pt_{\mathcal{N}_{i}}=\frac{1}{|\mathcal{N}_{i}|}\sum_{j}^{j\in\mathcal{N}_{i}}Sigmoid(w_{3}\times[u^{i}_{c},u^{n}_{i}])$ serves as the neighbors influences. Here $w_{2}$ and $w_{3}$ are randomly generated weight vectors that mimic the causal mechanism from the latent confounders to treatments. Then, $z_{i}$ can be directly obtained by network topology $E$ and $t_{\mathcal{N}_{i}}$.
We then modify the data generation of outcome $y$ in Jiang & Sun (2022):
$\small y_{i}(t_{i},z_{i})=t_{i}+z_{i}+po_{i}+0.5\times po_{\mathcal{N}_{i}}+e_{y,i},$
where $e_{y,i}$ is a Gaussian noise term, and $po_{i}=Sigmoid(w_{4}\times u_{i}+w_{5}\times u_{c})$, and $po_{\mathcal{N}_{i}}$ is the averages of $Sigmoid(w_{6}\times u_{c}+w_{7}\times u_{n})$. Here, $w_{4},w_{5},w_{6}$, and $w_{7}$ are all randomly generated weight vectors that mimic the causal mechanism from the confounders to outcomes. We denote this dataset as BC(homo) and Flickr(homo) [^1] Original datasets are available at https://github.com/songjiang0909/Causal-Inference-on-Networked-Data. since this generation of $y$ only measures the homogeneous causal effects.
Also following Chen et al. (2024), we consider the data generation of outcome $y$ with heterogeneous effect:
$\small y_{i}$ $(t_{i},z_{i})=t_{i}+z_{i}+po_{i}+0.5\times po_{\mathcal{N}_{i}}$
$+t_{i}(po_{i}+0.5\times po_{\mathcal{N}_{i}})+t_{i}(0.5\times po_{i}+po_{\mathcal{N}_{i}})+e_{y,i},$
which is denoted as BC(hete) and Flickr(hete).
Due to the space limit, we leave the detailed data generation process of the synthetic dataset in Appendix C.2.
Baselines We denote our method as Ours and a variant without IPM loss as Ours_w/o_IPM. We compare our methods with several state-of-the-art baselines  [^2] The details are in Appendix C.1. Our code will be available upon acceptance. Following Chen et al. (2024), we modify TARNET, CFR (Johansson et al., 2021) and ND (Guo et al., 2020) by additionally inputting the exposure $z_{i}$, denoted as TARNET+z, CFR+z and ND+z respectively. We also consider several baselines that are designed for networked effect estimation under the same setting, including GEst (Ma & Tresp, 2021), NetEst (Jiang & Sun, 2022), and TNet (Chen et al., 2024).
Metrics In this paper, we use the Mean Absolute Error (MAE) on AME, ASE, and ATE as our metric, i.e., $\varepsilon_{average}=|\hat{\tau}-\tau|$, where $\tau$ and $\hat{\tau}$ are the average causal effect and estimated one. We also use the Rooted Precision in Estimation of Heterogeneous Effect on IME, ASE, and ITE, $\varepsilon_{individual}=\sqrt{\frac{1}{n}\Sigma_{i=1}^{n}(\hat{\tau}_{i}-\tau_{i})^{2}}$, where $\tau_{i}$ and $\hat{\tau}_{i}$ are the individual causal effect and estimated one. The mean and standard deviation of these metrics via $5$ times running are reported. Note that our main estimands are AME, ASE, and ATE in this paper.
6.2 Experimental Analyses
Figure 4: Visualization of recovered and ground-true latent confounders $U^{i},U^{c}$, and $U^{n}$.
Effectiveness of Our Method As shown in Table 1 and Table 2, we have conducted experiments by running our proposed methods and several baselines. Overall, our methods outperform all methods consistently with smaller estimation errors in all metrics, indicating the effectiveness of our methods. Specifically, compared with the baselines, our methods perform better in terms of both average and heterogeneous treatment effect estimation, with smaller errors and standard deviation. This is reasonable since existing methods do not consider the latent confounders that hinder their identifications. Both Ours and Ours_w/o_IPM are based on identifiable representation learning techniques and thereby achieve superior performances with recovered latent confounders. Compared ours with its variant without the IPM term, we found that the IPM term can slightly improve the performance. This is due to the fact that the IPM term effectively mitigates the confounding bias with balanced representations.
$\hat{u}^{n}$ $\hat{u}^{c}$ $\hat{u}^{i}$
${u}^{i}$ 0.2489 0.2775 0.9505
${u}^{c}$ 0.2179 0.8930 0.5070
${u}^{n}$ 0.9435 0.4125 0.1412
Table 3: MCC results of recovered latent confounders.
Correctness of Representation Learning To validate the correctness of our representation learning method, We conduct experiments in the simulated dataset and visualize the recovered latent confounders $\hat{U}^{i},\hat{U}^{c},\hat{U}^{n}$ with ground-true $U^{i},U^{c},U^{n}$ in Figure 4. And we calculate the MCC results in Table 3. The result shows that the recovered latent confounders are highly correlated with the ground-true latent confounders, with very high MCC values. This indicates that our method can correctly recover the latent confounders, which validates the correctness of Theorem 4.2.
7 Conclusion
In this paper, we address the problem of networked causal effect identification and estimation in the presence of latent confounders. We leverage the networked information to achieve the identifiability of latent confounders. With identified latent confounders, we theoretically establish the identification result of networked effects. We further devise an effective estimator built on the theoretical findings. Extensive experiments validate the correctness of our theories and the effectiveness of our proposed estmator.
Impact Statement
This paper presents work whose goal is to advance the field of causal inference under networked interference. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Proof of Theorem 4.2
The proof can be directly obtained from the proof of Theorem 1 in Lu et al. (2022), with the following modifications:
  • we consider $X_{\mathcal{N}}$ as the prior conditional set;
  • we specify that the latent confounder $U$ contains three parts, i.e., $U=[U^{i},U^{c},U^{n}]$.
Appendix B Proof of Theorem 4.3
Proof.
Under Theorem 4.2, we can recover the whole distribution $p(x,u^{i},u^{c},u^{n},t,z,y)$, then the conditional average dose-response function is identified by
$\mu(t,z,x,x_{\mathcal{N}})$ $=\mathbb{E}[Y(t,z)|X=x,X_{\mathcal{N}}=x_{\mathcal{N}}]$ (12)
$=\mathbb{E}[\mathbb{E}[Y(t,z)|X=x,X_{\mathcal{N}}=x_{\mathcal{N}},U^{i}=u^{i},U^{c}=u^{c},U^{c}_{\mathcal{N}}=u^{c}_{\mathcal{N}},U^{n}_{\mathcal{N}}=u^{n}_{\mathcal{N}}]]$
$=\mathbb{E}[\mathbb{E}[Y(t,z)|X=x,X_{\mathcal{N}}=x_{\mathcal{N}},U^{i}=u^{i},U^{c}=u^{c},U^{c}_{\mathcal{N}}=u^{c}_{\mathcal{N}},U^{n}_{\mathcal{N}}=u^{n}_{\mathcal{N}},T=t,Z=z]]$
$=\mathbb{E}[\mathbb{E}[Y|X=x,X_{\mathcal{N}}=x_{\mathcal{N}},U^{i}=u^{i},U^{c}=u^{c},U^{c}_{\mathcal{N}}=u^{c}_{\mathcal{N}},U^{n}_{\mathcal{N}}=u^{n}_{\mathcal{N}},T=t,Z=z]]$
where the third equality is based on Assumption 3.11 and the forth equality is based on Assumption 3.7. and then the networked effect $\psi(t,z):=\mathbb{E}[Y(t,z)]$ is immediately identified. ∎
Appendix C Additional Experimental Details
C.1 Baseline Methods
The compared baselines in this paper are
  • TARNET+z: Original Tarnet (Johansson et al., 2021) uses two-heads neural networks, serving as T-Learner-like estimator, to estimate causal effects under no interference assumption. We modify TARNET by additionally inputting the exposure $z_{i}$.
  • CFR+z: Original CFR (Johansson et al., 2021) uses two-heads neural networks with an MMD term to achieve counterfactual regression under no interference assumption. We modify CFR by additionally inputting the exposure $z_{i}$.
  • ND+z: Original ND (Guo et al., 2020) propose network deconfounder framework by using network information under no interference assumption. We modify ND by additionally inputting the exposure $z_{i}$.
  • GEst (Ma & Tresp, 2021): GEst, based on CFR, uses GCN to aggregate the features of neighbors and input the exposure $z_{i}$ to estimate causal effects under networked interference.
  • NetEst (Jiang & Sun, 2022): NetEst learns balanced representation via adversarial learning for networked causal effect estimation.
  • TNet (Chen et al., 2024): TNet utilizes targeted learning techniques into its neural networks model to estimate causal effects under networked interference in a double robust manner.
C.2 Detailed Data Generation of Synthetic Dataset
We first generate the network graph $E$ with expected degrees $5$.
According to $E$, we directly generate all samples’ $u^{i},u^{c},u^{n}$ from a multivariate Gaussian distribution. We set the sample size as $1000$, then we first sample $\mathbf{u}=[[u^{i}]^{T},[u^{c}]^{T},[u^{n}]^{T}]^{T}$ as a $3000$-dimensional vector. To mimic the correlation among $u^{i},u^{c},u^{n}$ and between units, we generate the $3000\times 3000$ covariance matrix $V$ with all elements equal to $0.1$ besides diagonal equal to $1$. This results in that, for the same units, $u^{i}_{i},u^{c}_{i},u^{n}_{i}$ are correlated with covariance $0.1$, and for the different units, their latent confounders are also correlated with covariance $0.1$, e.g. $u_{i}^{i}$ and $u_{i}^{c}$. Then we generate $\mathbf{u}\sim\mathcal{N}(\mathbf{1},V)$ and reshape $\mathbf{u}$ as $[u^{i},u^{c},u^{n}]$, i.e., reshape to obtain the $1000\times 3$ design matrix of $u$.
Then, given $[u^{i},u^{c},u^{n}]$, we generate treatment $t$ as
$t_{i}\sim\textbf{Bern}(1/(1+\exp(-\frac{u_{i}^{i}+u_{i}^{c}+\sum_{j}^{j\in\mathcal{N}_{i}}(1.5u_{j}^{c}-0.5u_{j}^{n})-C}{4})))$ (13)
where $C$ is a constant set to ensure the positivity assumption holds. And then, $z$ can be obtained by aggregating $t$ using $E$.
Sequentially, the outcome $y$ is generated by
$y_{i}=u^{i}_{i}+u^{c}_{i}+t_{i}\times(2u^{i}+1.6u^{c})+z_{i}\times(\sum_{j}^{j\in\mathcal{N}_{i}}(1.5u_{j}^{c}-0.5u_{j}^{n})+1.5u^{i}_{i}+0.5u^{c}_{i})+\epsilon_{y,i},$ (14)
where $\epsilon_{y,i}$ is the Gaussian noise.
Finally, we generate the proxies $x$ of latent confounders by $x_{i}=w[u^{i}_{i},u^{c}_{i},u^{n}_{i}]+\epsilon_{x,i}$ where $w$ is a randomly $3\times 6$ weight matrix sampled from Gaussian distribution with a mean $1$ and scale $1$, and $\epsilon_{x,i}$ is the Gaussian noise. This results in the final $6$-dimensional observed proxies $x$.
C.3 Additional Experimental Results
We report additional experimental results on BC(homo) and BC(hete) datasets in Table 4 and Table 5. The results are similar to the experimental results in the Flickr dataset.
Table 4: Experimental results on BC(homo) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
$\varepsilon_{average}$ $\varepsilon_{individual}$
Within Sample Out-of Sample Within Sample Out-of Sample
Methods AME ASE ATE AME ASE ATE IME ISE ITE IME ISE ITE
TARNET+z $0.1573_{\pm 0.0405}$ $0.0824_{\pm 0.0149}$ $0.2046_{\pm 0.0193}$ $0.1492_{\pm 0.0370}$ $0.0855_{\pm 0.0147}$ $0.1982_{\pm 0.0380}$ $0.2096_{\pm 0.0250}$ $0.1161_{\pm 0.0159}$ $0.2444_{\pm 0.0239}$ $0.2809_{\pm 0.0507}$ $0.1209_{\pm 0.0152}$ $0.3062_{\pm 0.0627}$
CFR+z $0.0788_{\pm 0.0096}$ $0.1157_{\pm 0.0076}$ $0.2323_{\pm 0.0106}$ $0.0770_{\pm 0.0099}$ $0.1157_{\pm 0.0075}$ $0.2306_{\pm 0.0106}$ $0.0796_{\pm 0.0091}$ $0.1158_{\pm 0.0076}$ $0.2325_{\pm 0.0106}$ $0.1000_{\pm 0.0388}$ $0.1157_{\pm 0.0075}$ $0.2405_{\pm 0.0166}$
GEst $0.1872_{\pm 0.0672}$ $0.2369_{\pm 0.0607}$ $0.1422_{\pm 0.0562}$ $0.1955_{\pm 0.0611}$ $0.2391_{\pm 0.0617}$ $0.1302_{\pm 0.0524}$ $0.2307_{\pm 0.0493}$ $0.2603_{\pm 0.0586}$ $0.1877_{\pm 0.0495}$ $0.2388_{\pm 0.0431}$ $0.2623_{\pm 0.0592}$ $0.1790_{\pm 0.0440}$
ND+z $0.2375_{\pm 0.0450}$ $0.0316_{\pm 0.0104}$ $0.0790_{\pm 0.0226}$ $0.2380_{\pm 0.0458}$ $0.0323_{\pm 0.0122}$ $0.0768_{\pm 0.0254}$ $0.2377_{\pm 0.0448}$ $0.0321_{\pm 0.0101}$ $0.0792_{\pm 0.0226}$ $0.2477_{\pm 0.0460}$ $0.0379_{\pm 0.0099}$ $0.1068_{\pm 0.0172}$
NetEst $0.1059_{\pm 0.0609}$ $0.0284_{\pm 0.0297}$ $0.0387_{\pm 0.0288}$ $0.0987_{\pm 0.0663}$ $0.0257_{\pm 0.0276}$ $0.0356_{\pm 0.0268}$ $0.1366_{\pm 0.0481}$ $0.0631_{\pm 0.0205}$ $0.0994_{\pm 0.0214}$ $0.1680_{\pm 0.0620}$ $0.0920_{\pm 0.0316}$ $0.1507_{\pm 0.0647}$
TNet $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $0.0473_{\pm 0.0229}$ $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $0.0473_{\pm 0.0229}$ $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $0.0473_{\pm 0.0229}$ $0.1045_{\pm 0.0610}$ $0.0502_{\pm 0.0559}$ $0.0473_{\pm 0.0229}$
Ours_w/o_IPM $0.0356_{\pm 0.0176}$ $0.0222_{\pm 0.0104}$ $0.0514_{\pm 0.0163}$ $0.0416_{\pm 0.0176}$ $0.0206_{\pm 0.0100}$ $0.0433_{\pm 0.0136}$ $0.0411_{\pm 0.0147}$ $0.0244_{\pm 0.0122}$ $0.0554_{\pm 0.0182}$ $0.0435_{\pm 0.0169}$ $0.0218_{\pm 0.0106}$ $0.0453_{\pm 0.0137}$
Ours $0.0661_{\pm 0.0485}$ $0.0232_{\pm 0.0164}$ $0.0442_{\pm 0.0258}$ $0.0661_{\pm 0.0485}$ $0.0232_{\pm 0.0164}$ $0.0442_{\pm 0.0258}$ $0.0661_{\pm 0.0485}$ $0.0232_{\pm 0.0164}$ $0.0442_{\pm 0.0258}$ $0.0661_{\pm 0.0485}$ $0.0232_{\pm 0.0164}$ $0.0442_{\pm 0.0258}$
Table 5: Experimental results on BC(hete) Dataset. The top result is highlighted in bold, and the runner-up is underlined.
$\varepsilon_{average}$ $\varepsilon_{individual}$
Within Sample Out-of Sample Within Sample Out-of Sample
Methods AME ASE ATE AME ASE ATE IME ISE ITE IME ISE ITE
TARNET+z $0.2538_{\pm 0.1127}$ $0.1657_{\pm 0.0563}$ $0.3866_{\pm 0.0711}$ $0.2619_{\pm 0.1054}$ $0.1701_{\pm 0.0594}$ $0.4044_{\pm 0.1334}$ $0.3455_{\pm 0.0654}$ $0.2122_{\pm 0.0558}$ $0.4605_{\pm 0.0720}$ $0.5590_{\pm 0.2643}$ $0.2207_{\pm 0.0510}$ $0.6349_{\pm 0.3280}$
CFR+z $0.1580_{\pm 0.0189}$ $0.2071_{\pm 0.0237}$ $0.4067_{\pm 0.0407}$ $0.1559_{\pm 0.0203}$ $0.2076_{\pm 0.0245}$ $0.4061_{\pm 0.0449}$ $0.1825_{\pm 0.0129}$ $0.2092_{\pm 0.0233}$ $0.4316_{\pm 0.0351}$ $0.2058_{\pm 0.0432}$ $0.2098_{\pm 0.0242}$ $0.4422_{\pm 0.0456}$
GEst $0.2734_{\pm 0.1240}$ $0.4257_{\pm 0.0973}$ $0.2916_{\pm 0.1119}$ $0.2722_{\pm 0.1308}$ $0.4277_{\pm 0.1012}$ $0.2873_{\pm 0.1220}$ $0.3334_{\pm 0.1082}$ $0.4592_{\pm 0.0934}$ $0.3546_{\pm 0.0947}$ $0.3832_{\pm 0.1474}$ $0.4612_{\pm 0.0972}$ $0.3958_{\pm 0.1486}$
ND+z $0.4124_{\pm 0.0702}$ $0.0451_{\pm 0.0201}$ $0.1330_{\pm 0.0205}$ $0.4111_{\pm 0.0737}$ $0.0486_{\pm 0.0206}$ $0.1326_{\pm 0.0261}$ $0.4226_{\pm 0.0673}$ $0.0562_{\pm 0.0146}$ $0.1941_{\pm 0.0246}$ $0.4330_{\pm 0.0662}$ $0.0666_{\pm 0.0137}$ $0.2211_{\pm 0.0321}$
NetEst $0.1643_{\pm 0.1337}$ $0.0450_{\pm 0.0180}$ $0.0594_{\pm 0.0262}$ $0.1857_{\pm 0.1168}$ $0.0405_{\pm 0.0252}$ $0.0343_{\pm 0.0179}$ $0.2199_{\pm 0.1039}$ $0.0667_{\pm 0.0171}$ $0.1731_{\pm 0.0368}$ $1.5595_{\pm 2.5329}$ $1.1347_{\pm 1.9028}$ $1.0924_{\pm 1.7278}$
TNet $0.1216_{\pm 0.0864}$ $0.0537_{\pm 0.0524}$ $0.0429_{\pm 0.0301}$ $0.1257_{\pm 0.0727}$ $0.0537_{\pm 0.0511}$ $0.0481_{\pm 0.0269}$ $0.1731_{\pm 0.0450}$ $0.0655_{\pm 0.0465}$ $0.1458_{\pm 0.0175}$ $0.1915_{\pm 0.0542}$ $0.0650_{\pm 0.0458}$ $0.1740_{\pm 0.0621}$
Ours_w/o_IPM $0.0706_{\pm 0.0324}$ $0.0278_{\pm 0.0079}$ $0.0605_{\pm 0.0504}$ $0.0755_{\pm 0.0420}$ $0.0294_{\pm 0.0052}$ $0.0664_{\pm 0.0457}$ $0.1240_{\pm 0.0237}$ $0.0491_{\pm 0.0060}$ $0.1706_{\pm 0.0311}$ $0.1273_{\pm 0.0276}$ $0.0481_{\pm 0.0054}$ $0.1705_{\pm 0.0269}$
Ours $0.0625_{\pm 0.0597}$ $0.0195_{\pm 0.0162}$ $0.0579_{\pm 0.0255}$ $0.0604_{\pm 0.0610}$ $0.0195_{\pm 0.0130}$ $0.0598_{\pm 0.0259}$ $0.1175_{\pm 0.0380}$ $0.0373_{\pm 0.0099}$ $0.1545_{\pm 0.0141}$ $0.1172_{\pm 0.0392}$ $0.0365_{\pm 0.0074}$ $0.1563_{\pm 0.0130}$